15m-9-6m=9 I need someone to show me step by step how to work out this problem
Found 2 solutions by checkley71, mathdumb:
Answer by checkley71(8403)

COMBINE THE M's AND THE DIGITS THUS 15M-9-6M=9 OR 15M-6M=9+9 OR 9M=18 OR M=18/9 OR M=2

Answer by mathdumb(3)

15m-9-6m=9
15m-6m-9=9
9m-9=9
9m-9+9=9+9
9m=18
m=18/9
m=2